Five years hence, the age of Jacob will be three times that of his son. Five years ago, Jacob’s age was seven times that of his son. What are their present ages?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student
Here is the solution:

Let the age of Jacob and his son be x and y respectively.
According to the question,
(x + 5) = 3(y + 5) x – 3y = 10 …………………………………….. (1)
(x – 5) = 7(y – 5) x – 7y = -30 …………………………………….... (2)
From (1), we get
x = 3y + 10 ……………………. (3)
Substituting the value of x in (2),
we get
3y + 10 – 7y = -30 -4y = -40 y = 10 ………………… (4)
Substituting the value of y in (3),
we get
x = 3 x 10 + 10 = 40
Hence, the present age of Jacob’s and his son is 40 years and 10 years respectively.

Thanks